WAC Notes July 23 2019

From PREX Wiki
Jump to: navigation, search
Documentation
HOW TOs for shift crew
Expert Tools
All Expert Contacts



PREX Main << Weekly Analysis Coordinator << WAC Notes << Instructions

WAC Notes July 22 2019 << >> WAC Notes July 24 2019

July 23 2019

  • Slug 12 is the first without the old PREX 1 quartz (in LHRS) - from here on we are using different quartz: HALOG

Production Run List

Start Slug 12 at junk run 3583, IHWP In, FFB on always, no RHRS, new PREX II quartz (in LHRS)

Non-Production Runs and Bad Runs

  • Junk 3583: adaq1 test at midnight (slug12 starts here)
  • Junk 3584-  : Test runs after IHWP in

Pedestal/Calibrations Updates

  • Updated prex_maindet.3583-.map
    • Removed not_blindable and not_normalizable from RHRS MAIN detectors
    • Removed not_normalizable from RHRS AT detectors
  • Updated prexCH_beamline_eventcuts.3583-.map for 70uA running
    • bcm, bcm_an_ds, 55, 1e6, g, 0.8
    • bcm, bcm_an_us, 55, 1e6, g, 0
    • all other cuts as of before
  • Based on HALOG
    • Updated prexCH_beamline_pedestal.3583-.map
    • Updated prex_maindet_pedestal.3583-.map
    • Updated prex_sam_pedestal.3583-.map
      • Pedestal map files need to be updated once we have results from new current scan

Issues

Other Status Updates

  • Accumulated charge =
  • Slug 12 is the first without the old PREX 1 quartz in LHRS (RHRS still has old PREX 1 quartz)- from here on we are using different quartz: HALOG
  • I was informed that the prexPrompt_pass1 root file for run 3425 was accidentally overwritten
    • Re-prompted this run
  • Re-prompted run3571 ... the glitch is fixed, plots look ok!. The blinder is still 0'd out
  • Re-prompted run3573 ... event cut applied (CodaEventNumber>132000&&CodaEventNumber<172000).